at XX e random variables endh withy the Also, the correlation coeficients of Xi, x2, x3 are ρ12 =-0.3,P13 Find the mean and variance of 0.6, p23 = 0.5
Answer #1

E(X)-1%(2-1)dr-5 За (Decima

E(X2)-1%(2-r) dr-20 За (Decimal 0.45)

Var(X) = E(X^2) - (E(X))^2

= 0.45 - 0.625^2

= 0.059375

E(Y) = E(X1 + 3 X2 + 2 X3) = E(X) + 3 E(X) + 2 E(X)

= 6 E(X) = 6 *5/8 = 3.75

Var(Y) = Var(X1 + 3 X2 + 2 X3)= Var(X) + 9 Var(X) + 4 Var(X) + 2 * 3 *P12 Var(X) + 2 * 2 *ho_{13} Var(X) +2*)23*2*3*Var(X)

= Var(X) (14 + 6 * P12 + 4*ho_{13} + 12 )23)

= 0.059375* ( 14 + 6 * (-0.3) + 4 * 0.6 + 12 * 0.5)

= 1.223125